One person can do a certain job in twenty-one minutes, and another person can do the same job in twenty-eight minutes. How many minutes will it take them to do the job altogether?
Answer by checkley71(8403)

21*28/(21+28)=588/49=12 MINUTES TO COMPLETE THE JOB WORKING TOGETHER.